I NEED STEP BY STEP SOLUTIONS FOR EACH PART. DO NOT SKIP ANY PART. ANSWERS TO EACH ARE MARKED IN BOLD

   14. Suppose your venture’s expected mean cash flows are $(85,000) initially, followed by expected mean cash flows at the end of the first, second, and third years of $40,000, $40,000, and $35,000. What is the internal rate of return?

                        a. 13.9%

                        b. 14.7%

                        c. 16.2%

                        d. 17.2%

                        e. 19.2%        

16. Estimate the value of a privately-held firm based on the following information: stock price of a comparable firm = $20.00; net income of a comparable firm = $20,000; number of shares outstanding for the comparable firm = 10,000; and earnings per share for the target firm = $3.00.

                        a.   $10.00      

                        b.   $20.00

                        c.   $30.00

                        d.   $40.00

                        e.   $50.00

17. Estimate the value of a privately-held firm based on the following information: total market value (or capitalization value) of a comparable firm = $200,000; net income of a comparable firm = $40,000; number of shares outstanding for the comparable firm = 20,000; net income for the target firm = $15,000; and number of shares outstanding for the target firm = 10,000.

                        a.   $5.00        

                        b.   $7.50

                        c.   $10.00

                        d.   $12.50

                        e.   $15.00
Answer #1

14. IRR = where NPV (Net present value) is equal to zero

Year Cash flow PVF (1796) PV (1796) PVF (1896) PV (1896) 085000 1 40000 0.854701 34188.03 0.847458 33898.31 2 40000 0.730514

Using Trial & error approch we get

IRR = 17 % + ( 261.5457 - 0) / [ 261.5457 - (-1072.24)]

= 17% + 0.196

= 17.20 %

16. Stock price of a comparable firm = $20

Net income of a comparable firm = $20,000

Number of shares outstanding for the comparable firm = 10,000;

Earning per share o fcomparable firm = 20000 /10000 = $2

Earnings per share for the target firm = $3.00.

Value of Privately held firm

= ($20/$2) * $3

= $30

17. Net income for the target firm = $15,000; and n

Number of shares outstanding for the target firm = 10,000.

Earning per share of target firm = $ 15000 / 10000 = 1.5 per share

Total market value (or capitalization value) of a comparable firm = $200,000; Net income of a comparable firm = $40,000

M = 200000 / 40000 = 5

Value of privately held firm = 1.5 * 5

= 7.5